OpenStudy (anonymous):

Thermodynamics :

OpenStudy (anonymous):

When the system goes from the state A to the state B, through the trajectory A-C-B, absorbs a heat of 20,000 calories and makes a work of 7,500 calories. a) How much heat absorbs the system through the trajectory A-D-B, if the work done is 2,500 calories b) When the system returns to the state A from the state B, through the curve, the work done is 5,000 calories. How much heat is absorbed or released? c) If Ua = 0 cal, and Ud = 10,000 cal, calculate the heat exchanged in the processes A-D and D-B. (U : internal energy) Answers : a) Q =15,000 cal ; b) Q= -17,500 cal ; c) Qad= 12,500 cal ; d) Qdb = 2,500 cal

OpenStudy (anonymous):

Look at the cycle in the attachment

OpenStudy (nincompoop):

everytime a work is done the energy is negative, and when absorb then it's positive. so what you have is work + absorb

OpenStudy (anonymous):

I think in engineering, we assume that the work done is positive

OpenStudy (nincompoop):

no. for every time a work is done, you are releasing energy, so that is a negative value. the engineering concept is not different from physics or chemistry.

OpenStudy (nincompoop):

the only time it becomes tricky is when you are dealing with living systems, since living things cannot use heat as energy.

OpenStudy (anonymous):

Then, I have to use U = Q - W ???

OpenStudy (nincompoop):

ya

OpenStudy (anonymous):

Ah, I MADE IT !!

OpenStudy (nincompoop):

first law of thermodynamics

OpenStudy (nincompoop):

you know, even if you use w + q, you will still get the same answer. that is to make sure that w is negative :D

OpenStudy (nincompoop):

look at the problem when: \[\Delta U=Q-W\] if W is negative then your equation changes to \[\Delta U=Q-(-W)=Q+W\] it becomes mathematically incoherent, yes?

OpenStudy (anonymous):

THe internal energy variation is equal in the process A-C-B and the process A-D-B \[\Delta Ucb = Qcb - 7,500 \cal\] \[\Delta Uac = Qac - 0 (isocoric) \] \[Qcb = \Delta Ucb + 7,500 \cal\] \[Qac + Qcb = 20,000 \cal\] \[\Delta Uac + \Delta Ucb + 7,500 \cal = 20,000 \cal \] \[\Delta Uac + \Delta Ucb = 12,500 calories\] \[\Delta Uad + \Delta Udb = 12,500 [calories]\] \[Qad - 2,500 + Qdb = 12,500 \] \[Qad + Qdb = 15,000 [calories]\]

OpenStudy (nincompoop):

I didn't notice the graph LOL!!!

OpenStudy (anonymous):

Hahahahah. I didn't remember that when a process is reversible, it only matters the initial and final states, so the internal energy variation in both ways is the same